OpenStudy (rayford):

Solve x2 + 9x + 9 = 0. using quadratic formula

mathslover (mathslover):

So here is a hint : \[\huge{x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}}\]

OpenStudy (king):

where a,b,c are coeffiecients of a quadratic eqn. ax^2+bx+c

mathslover (mathslover):

x is root of the quadratic equation : \(\huge{ax^2+bx+c=0}\)

OpenStudy (king):

substitute a,b,c in the formula to get x u will get 2 values as there are 2 roots of a quad. eqn.

OpenStudy (rayford):

is it -9root117 all over 2?

OpenStudy (rayford):

guys i dont get this man..all of my answer choices have a 10 at the end

OpenStudy (rayford):

oops sorry wrong question.

OpenStudy (ash2326):

@rayford you got it?

OpenStudy (rayford):

is it -9 + or minus root 117 all over 2?

OpenStudy (ash2326):

check again, there is a small mistake Check the term in square root

OpenStudy (rayford):

3root5?

OpenStudy (ash2326):

Correct, it'd be \[\frac{-9\pm 3\sqrt 5}{2}\]

OpenStudy (rayford):

YES!!!!!! 100 ON MY EXAM YEAS!! ty i understand now :[)

OpenStudy (ash2326):

@rayford were you giving a test?

OpenStudy (rayford):

oh no! i was stating that 100 on my exam here i come. cause now that i understand i will do good on my exam :]

OpenStudy (ash2326):

Great:D wish you great luck with that. Let me know whenever you need help here. :D

OpenStudy (rayford):

thank you very much. lol champion